### v1

@@ 5 Concluding remarks
This paper develops a tractable model of strategic manipulation by artists on music streaming platforms under the pro-rata revenue sharing rule. The framework is flexible enough to accommodate evolving detection methods and increasingly sophisticated fraud strategies, including those enabled by machine learning and advanced identification technologies. It shows that pro-rata can be robust to fraud when fraud technology is weak, yielding an efficient fraud-free equilibrium. When fraud technology is strong, a unique fraud equilibrium exists, but aggregate fake streams remain bounded because increased total streams dilute the value of manipulation. Fraud may also compress payoffs and improve egalitarian fairness despite efficiency losses.
@@ A.1 Proof of Lemma 1
$v_{i}(t_{i},\tau_{i})=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V-t_{i}+d_{i}.$ (5)
To show $t_{i}=0$ is a strict dominant strategy of agent $i$, one needs to show that for all $\tau_{i}\geq 0$, it holds for all $t_{i}>0$ that $v_{i}(0,\tau_{i})>v_{i}(t_{i},\tau_{i})$. It is
@@
To sum up, if $d_{i}\leq\frac{1}{2}$, $V\leq\frac{1}{1-d_{i}}$; if $d_{i}>\frac{1}{2}$, $V\leq 4d_{i}$.
@@ A.2 Proof of Lemma 2
$\frac{\partial v_{i}(t_{i},\tau_{j})}{\partial t_{i}}=\frac{1-d_{i}+\tau_{i}}{(1+t_{i}+\tau_{i})^{2}}V-1,$ (9)
@@
$\frac{\partial^{2}v_{i}(t_{i},\tau_{j})}{\partial t_{i}^{2}}=-2\frac{(1-d_{i}+\tau_{i})}{(1+t_{i}+\tau_{i})^{3}}V,$
@@
To sum up, if $V>\frac{1}{1-d_{\min}}$, there exists a unique fraud equilibrium, where at least one agent takes positive strategy.
@@ A.3 Proof of Theorem 2
Proof. (a) Given $i\in N^{d}$ and $d_{j}\leq d_{i}$, suppose $j\in N\setminus N^{d}$. There is $t^{*}_{i}>0$, $t^{*}_{j}=0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}=0$, and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}\leq 0$. From (10), it implies
@@
(b) Conversely, given $i\in N\setminus N^{d}$ and $d_{j}\geq d_{i}$, suppose $j\in N^{d}$. There is $t^{*}_{i}=0$, $t^{*}_{j}>0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}\leq 0$ and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}=0$. From (10), it implies
@@
which is a contradiction. Thus $j\in N\setminus N^{d}$.
@@ A.4 Proof of Proposition 1
$\overline{r}^{\prime}(s)$ $=\frac{1}{2s^{2}}V\Bigg[\frac{\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}+1\Bigg].$
$=\frac{1}{2s^{2}}V\frac{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}+\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}$
$=\frac{1}{2s^{2}}V\frac{\overline{r}(s)-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}},$
@@
which approaches its upper bound $V$ as $n\rightarrow+\infty$. Combining with the fact that $\underline{r}(n)$ also approaches $V$ as well. The upper bound of $1+T^{*}$ is thus $V$.
@@ A.5 Proofs of Corollaries 1, 2, and 3
Proof of Corollary 1. From statement (a) of Theorem 2, equilibrium strategy of agent $i\in N^{d}$ is decreasing in $d_{i}$. All agents taking positive strategies is thus equivalent to that the agent with $d_{\mathrm{max}}$ takes positive strategy. Suppose agent $i$ is the agent with $d_{i}=d_{\mathrm{max}}$. From (10), the equilibrium strategy of agent $j$ is
$d_{\mathrm{max}}+t_{j}^{*}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$
Then $t_{j}^{*}>0$ is equivalent to
@@
which is exactly $d_{j}\geq d^{*}$. Substituting (13) into (15), formula shown in corollary can be derived.
@@
which is equivalent to the formula.
@@ A.6 Proof of Proposition 2
Consider two utility profiles $u^{*}$ and $u^{0}$, corresponding to that in fraud equilibrium and the fraud-free situation. Obviously, $u_{i}^{0}=u_{i}(\mathbf{0})=d_{i}(1+V)$ for all $i\in N$. From (10) and (15), for all agent $i\in N^{d}$ with positive equilibrium strategy, there is
@@
$u_{j}^{*}=d_{j}\bigg(1+\frac{V}{1+T^{*}}\bigg)=\frac{d_{j}}{1+T^{*}}V-d^{*}+d^{*}+d_{j}>\frac{d^{*}}{1+T^{*}}V-d^{*}+2d_{i}=u_{i}^{*}.$
Thus we can conclude that utilities in the fraud equilibrium follow the same order as those in the fraud-free situation.
To show the fairness property, the comparison between two profiles for the utility of pivotal agent $i$ with $d_{i}=d_{\min}$ is needed. Combining (16) and $u_{i}^{0}=d_{\min}(1+V)$, there is
$u_{j}^{*}-u_{j}^{0}$ $=\frac{d^{*}}{1+T^{*}}V-d^{*}+2d_{\min}-d_{\min}(1+V)$
$=\frac{1}{V}[(1+T^{*})-V]^{2}-d_{\min}(V-1).$
So that the fraud equilibrium is fairer, i.e., $u_{i}^{*}>u_{i}^{0}$, is equivalent to
$[(1+T^{*})-V]^{2}>d_{\min}V(V-1).$
That is, either $1+T^{*}>V+\sqrt{d_{\min}V(V-1)}$, or $1+T^{*}<V-\sqrt{d_{\min}V(V-1)}$. From Proposition 1, there is $1+T^{*}\leq V$. Thus only the later range is kept.
@@ A.7 Proof of Theorem 3
$v_{i}(t_{i},\tau_{i};\alpha)=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V\alpha-t_{i}+d_{i}\alpha+\frac{\lambda_{0}}{n\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha)$ (17)
$t_{i}=0$ is a strict dominant strategy of agent $i$ is equivalent to that for all $\tau_{i}\geq 0$, $v_{i}(0,\tau_{i};\alpha)>v_{i}(t_{i},\tau_{i};\alpha)$ holds for all $t_{i}>0$. That is,
@@
which is exactly (6) if $V$ is replaced with $V\alpha$. Repeating the analysis of Lemma 1 gives the equivalent formula. That is, $V\alpha\leq\frac{1}{1-d_{i}}$ if $d_{i}\leq\frac{1}{2}$, and $V\alpha\leq 4d_{i}$ if $d_{i}>\frac{1}{2}$. Thus the theorem can be immediately derived.
A.8 Existence of equilibrium under slightly high threshold
@@
$u_{i}(\widehat{t})=\frac{\widehat{d}}{1+\widehat{T}}V-\widehat{d}+2d_{\min}\geq\frac{d^{*}+(\widehat{d}-d^{*})}{1+T^{*}+n(\widehat{d}-d^{*})}V-d^{*}-(\widehat{d}-d^{*})+2d_{\min}.$
@@
$H(\theta)=\frac{d^{*}+\theta}{1+T^{*}+n\theta}V-d^{*}-\theta+2d_{\min}.$
@@
From two cases, we have $H(\theta^{*})\geq H(0)=u_{i}(t^{*})>u_{i}(0,t^{*}_{-i})\geq 0$, and $H(V+d_{\min})=(\frac{d^{*}+V+d_{\min}}{1+T^{*}+nV+nd_{\min}}-1)V-(d^{*}-d_{\min})<0$. Thus there exists a $\theta^{**}\in(\theta^{*},V+d_{\min})$ such that $H(\theta^{**})=0$. Combining that $H(\theta)$ is strictly increasing on $(0,\theta^{*})$ for positive $\theta^{*}$, and strictly decreasing on $(\theta^{*},V+d_{\min})$, it has $H(\theta)\geq 0$ for all $\theta\in[0,\theta^{**}]$. Thus we can find a non-empty range for $\widehat{d}\in(d^{*},d^{*}+\theta^{**}]$, such that $u_{i}(\widehat{t})\geq H(\theta^{**})\geq 0$, verifying the existence of new equilibrium.

### v2

@@ 7 Concluding remarks
This paper develops a tractable model of strategic manipulation by artists on music streaming platforms under the pro-rata revenue sharing rule. The framework is flexible enough to accommodate evolving detection methods and increasingly sophisticated fraud strategies, including those enabled by machine learning and advanced identification technologies. It shows that pro-rata can be robust to fraud when fraud technology is weak, yielding an efficient fraud-free equilibrium. When fraud technology is strong, a unique fraud equilibrium exists, but aggregate fake streams remain bounded because increased total streams dilute the value of manipulation. Fraud may also compress payoffs and improve egalitarian fairness despite efficiency losses. Therefore, this paper defends the pro-rata revenue sharing rule against concerns about click-fraud.
@@ A.1 Proof of Lemma 1
$v_{i}(t_{i},\tau_{i})=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V-t_{i}+\frac{d_{i}}{\xi}.$ (6)
To show $t_{i}=0$ is a strictly dominant strategy of agent $i$, one needs to show that for all $\tau_{i}\geq 0$, it holds for all $t_{i}>0$ that $v_{i}(0,\tau_{i})>v_{i}(t_{i},\tau_{i})$. It is
@@
To sum up, if $d_{i}\leq\frac{1}{2}$, $V\leq\frac{1}{1-d_{i}}$; if $d_{i}>\frac{1}{2}$, $V\leq 4d_{i}$. ∎
@@ A.2 Proof of Lemma 2
$\frac{\partial v_{i}(t_{i},\tau_{i})}{\partial t_{i}}=\frac{1-d_{i}+\tau_{i}}{(1+t_{i}+\tau_{i})^{2}}V-1,$ (10)
@@
$\frac{\partial^{2}v_{i}(t_{i},\tau_{i})}{\partial t_{i}^{2}}=-2\frac{(1-d_{i}+\tau_{i})}{(1+t_{i}+\tau_{i})^{3}}V,$
@@
To sum up, if $V>\frac{1}{1-d_{\min}}$, there exists a unique fraud equilibrium, where at least one agent takes positive strategy. ∎
@@ A.3 Proof of Theorem 2
Proof. (a) Given $i\in N^{d}$ and $d_{j}\leq d_{i}$, suppose $j\in N\setminus N^{d}$. There is $t^{*}_{i}>0$, $t^{*}_{j}=0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}=0$, and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}\leq 0$. From (11) and (12), it implies
@@
(b) Conversely, given $i\in N\setminus N^{d}$ and $d_{j}\geq d_{i}$, suppose $j\in N^{d}$. There is $t^{*}_{i}=0$, $t^{*}_{j}>0$, $\frac{\partial v_{i}(t_{i}^{*},\tau_{i}^{*})}{\partial t_{i}}\leq 0$ and $\frac{\partial v_{j}(t_{j}^{*},\tau_{j}^{*})}{\partial t_{j}}=0$. From (11) and (12), it implies
@@
which is a contradiction. Thus $j\in N\setminus N^{d}$. ∎
@@ A.4 Proof of Proposition 1
$\begin{split}\overline{r}^{\prime}(s)&=\frac{1}{2s^{2}}V\Bigg[\frac{\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}+1\Bigg].\\ &=\frac{1}{2s^{2}}V\frac{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}+\frac{s-1}{2s}V-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}}\\ &=\frac{1}{2s^{2}}V\frac{\overline{r}(s)-1}{\sqrt{(\frac{s-1}{2s}V)^{2}+\frac{1}{s}V}},\end{split}$
@@
which approaches its upper bound $V$ as $n\rightarrow+\infty$. Combining with the fact that $\underline{r}(n)$ also approaches $V$ as well. The upper bound of $1+T^{*}$ is thus $V$. ∎
@@ A.5 Proofs of Corollaries 1, 2, and 3
Proof of Corollary 1. From statement (a) of Theorem 2, equilibrium strategy of agent $i\in N^{d}$ is decreasing in $d_{i}$. All agents taking positive strategies is thus equivalent to that the agent with $d_{\mathrm{max}}$ takes positive strategy. Suppose agent $i$ is the agent with $d_{i}=d_{\mathrm{max}}$. From (11), the equilibrium strategy of agent $i$ is
$d_{\mathrm{max}}+t_{i}^{*}=1+T^{*}-\frac{1}{V}(1+T^{*})^{2}.$
Then $t_{i}^{*}>0$ is equivalent to
@@
which is exactly $d_{j}\geq d^{*}$. Substituting (14) into (16), formula shown in corollary can be derived. ∎
@@
which is equivalent to the formula. ∎
@@ A.6 Proof of Proposition 2
Consider two utility profiles $u^{*}$ and $u^{0}$, corresponding to that in fraud equilibrium and the fraud-free situation. Obviously, $u_{i}^{0}=u_{i}(\mathbf{0})=d_{i}(\frac{1}{\xi}+V)$ for all $i\in N$. From (11) and (16), for all agent $i\in N^{d}$ with positive equilibrium strategy, there is
@@
$u_{j}^{*}=d_{j}\bigg(\frac{1}{\xi}+\frac{V}{1+T^{*}}\bigg)=\frac{d_{j}}{1+T^{*}}V-d^{*}+d^{*}+\frac{1}{\xi}d_{j}>\frac{d^{*}}{1+T^{*}}V-d^{*}+d_{i}+\frac{1}{\xi}d_{i}=u_{i}^{*},$
where the inequality is obtained from $d_{j}\geq d^{*}>d_{i}$. Thus we can conclude that utilities in the fraud equilibrium follow the same order as those in the fraud-free situation.
To show the fairness property, the comparison between two profiles for the utility of pivotal agent with $d_{\min}$ is needed. Denote pivotal agent’s utility in fraud equilibrium and that in fraud-free situation by $u^{*}_{\min}$ and $u^{0}_{\min}$, respectively. Combining (17) and $u_{i}^{0}=d_{i}(\frac{1}{\xi}+V)$, there is
$\begin{split}u^{*}_{\min}-u^{0}_{\min}&=\frac{d^{*}}{1+T^{*}}V-d^{*}+\bigg(1+\frac{1}{\xi}\bigg)d_{\min}-d_{\min}\bigg(\frac{1}{\xi}+V\bigg)\\ &=\frac{1}{V}[(1+T^{*})-V]^{2}-d_{\min}(V-1).\end{split}$
So that the fraud equilibrium is fairer, i.e., $u^{*}_{\min}>u^{0}_{\min}$, is equivalent to
$(1+T^{*})-V]^{2}>d_{\min}V(V-1).$
That is, either $1+T^{*}>V+\sqrt{d_{\min}V(V-1)}$, or $1+T^{*}<V-\sqrt{d_{\min}V(V-1)}$. From Proposition 1, there is $1+T^{*}\leq V$. Thus only the later range is kept. ∎
@@ A.7 Proof of Theorem 3
$v_{i}(t_{i},\tau_{i};\alpha)=\frac{d_{i}+t_{i}}{1+t_{i}+\tau_{i}}V\alpha-t_{i}+\frac{1}{\xi}\Bigg[d_{i}\alpha+\frac{\lambda_{0}}{m\overline{\lambda}}\sum_{k\in M}\pi_{ik}(1-\alpha)\Bigg].$ (18)
$t_{i}=0$ is a strictly dominant strategy of agent $i$ is equivalent to that for all $\tau_{i}\geq 0$, $v_{i}(0,\tau_{i};\alpha)>v_{i}(t_{i},\tau_{i};\alpha)$ holds for all $t_{i}>0$. That is,
@@
which is exactly (7) if $V$ is replaced with $V\alpha$. Repeating the analysis of Lemma 1 gives the equivalent formula. That is, $V\alpha\leq\frac{1}{1-d_{i}}$ if $d_{i}\leq\frac{1}{2}$, and $V\alpha\leq 4d_{i}$ if $d_{i}>\frac{1}{2}$. Thus the theorem can be immediately derived. ∎
A.8 Existence of slightly high threshold
@@
$\begin{split}u_{i}(\widehat{t})&=\frac{\widehat{d}}{1+\widehat{T}}V-\widehat{d}+\bigg(1+\frac{1}{\xi}\bigg)d_{\min}\\ &\geq\frac{d^{*}+(\widehat{d}-d^{*})}{1+T^{*}+n(\widehat{d}-d^{*})}V-d^{*}-(\widehat{d}-d^{*})+\bigg(1+\frac{1}{\xi}\bigg)d_{\min}.\end{split}$
@@
$H(\theta)=\frac{d^{*}+\theta}{1+T^{*}+n\theta}V-d^{*}-\theta+\bigg(1+\frac{1}{\xi}\bigg)d_{\min}.$
@@
From two cases, we have $H(\theta^{*})\geq H(0)=u_{i}(t^{*})>u_{i}(0,t^{*}_{-i})\geq 0$, and $H(V+(1+\frac{1}{\xi})d_{\min})=[\frac{d^{*}+V+(1+\frac{1}{\xi})d_{\min}}{1+T^{*}+nV+n(1+\frac{1}{\xi})d_{\min}}-1]V-d^{*}<0$, where $d^{*}+V+(1+\frac{1}{\xi})d_{\min}<1+T^{*}+nV+n(1+\frac{1}{\xi})d_{\min}$. Thus there exists a $\theta^{**}\in(\theta^{*},V+(1+\frac{1}{\xi})d_{\min})$ such that $H(\theta^{**})=0$. Combining that $H(\theta)$ is strictly increasing on $(0,\theta^{*})$ for positive $\theta^{*}$, and strictly decreasing on $(\theta^{*},V+(1+\frac{1}{\xi})d_{\min})$, it has $H(\theta)\geq 0$ for all $\theta\in[0,\theta^{**}]$. Thus we can find a non-empty range for $\widehat{d}\in(d^{*},d^{*}+\theta^{**}]$, such that $u_{i}(\widehat{t})\geq H(\theta^{**})\geq 0$, verifying the existence of slightly high threshold for the new equilibrium. ∎